a 60.0kg shell is shot from a cannon to a height of 400.0m What is the potential energy at this height?How much will it's potential energy decrease when the shell falls to a height of 150m
1
Expert's answer
2018-11-08T11:18:09-0500

The gravitational potential energy

U(h)=mgh=60.0 kg×9.80 N/kg×400 m=2.35×〖10〗^5 J

U(H)=mgH=60.0 kg×9.80 N/kg×150 m=0.88×〖10〗^5 J

ΔU=U(h)-U(H)=1.47×〖10〗^5 J